# A Block-Shifted Cyclic Reduction Algorithm for Solving a Class of Quadratic Matrix Equations

### v1

@@ Abstract
organization=Lanzhou University of Technology,city=Lanzhou, country=China
@@
organization=University of Pisa,city=Pisa, country=Italy
@@ 1 Introduction
A commonly used numerically stable approach for solving QMEs is to linearize them into associated generalized eigenvalue problems, which can then be efficiently solved using the QZ algorithm to compute all the solutions [11, 8, 12]. However, the computational cost, limits its practicality for large-scale problems. Iterative methods, such as Newton’s method [11, 13], the Bernoulli iteration (BI) and its variants [11, 14], the fixed-point iterations [15, 16, 17, 18, 11, 3, 19], offer greater efficiency but may converge slowly or stagnate in nearly null-recurrent QBDs. An alternate to classical algorithms is the cyclic reduction (CR) algorithm. Originally developed by Golub and Hockney for block tridiagonal systems [20, 21], the CR algorithm was later extended to solve nonlinear matrix equations based on a functional representation [3, 22]. The central idea is to transform the nonlinear matrix equation into an equivalent semi-infinite, block Toeplitz, block tridiagonal or Hessenberg linear system and apply CR to solve this system. The comprehensive survey [23] provides a detailed account of the development and applications of CR.
@@ 2 Preliminaries and assumptions
Assume $a(z)$ not identically zero and denote its roots as $\lambda_{i}$, $i=1,\ldots,2m$, adding $q$ zeros at infinity if $a(z)$ has degree 2mq2m-q. Without loss of generality, assume that the roots are ordered according to their modulus as
@@ 2.1 Cyclic Reduction (CR)
Concerning convergence properties, we recall the following result, which follows from Theorems 9 and 10 of [23] (see also Theorem 7.6 in [3, Section 7.3]), and expresses the quadratic convergence:
@@
In the case where $\lambda_{m}=\lambda_{m+1}=1$, $|\lambda_{m-1}|<1$ and $|\lambda_{m+2}|>1$, under mild conditions, the convergence of CR turns to linear [29]. The convergence is restored to quadratic by applying a shifting technique to remove the eigenvalue 1 from the eigenvalues of $A(z)$ ([30], [3, Section 9.2]).
@@ 2.2 Assumptions
$A(z)=(I_{m}-zR)W(zI_{m}-G),$
where $W$ is a suitable nonsingular matrix; moreover, the matrix $A_{1}+A_{2}G$ is nosingular, and the matrices $G$ and $R$ are related by the equation (see [3, Section 3.3]):
@@
In this case, $\mu_{1},\ldots,\mu_{\ell}$ are the $\ell$-th roots of $1$. Moreover, $G$ and $R$ have nonnegative entries and, among all the possible solutions of (3) and (4), respectively, are the minimal ones according to the component-wise ordering (for details on these topics we refer the reader to [2], [3, Section 4.4], [35]).
@@ 2.3 Block shift-and-deflate technique
The following results from [33] show how a set of eigenvalues of $A(z)$ can be moved to 0 and to infinity.
@@
Let $1\leq q<m$, $S_{1}\in\mathbb{C}^{q\times q}$ with $\det S_{1}\neq 0$, $S_{2}\in\mathbb{C}^{q\times q}$, and $U_{1}\in\mathbb{C}^{q\times m}$, $V_{2}\in\mathbb{C}^{m\times q}$ be full rank matrices such that
$S_{1}^{2}U_{1}A_{0}+S_{1}U_{1}A_{1}+S_{2}A_{2}=0,\penalty 10000\ \penalty 10000\ A_{0}V_{2}+A_{1}V_{2}S_{2}+A_{2}V_{2}S_{2}^{2}=0.$
@@
$\widetilde{A}(z)=\left(I_{m}+zX(S_{1}-zI_{q})^{-1}U_{1}\right)A(z)\left(I_{m}+V_{2}(zI_{q}-S_{2})^{-1}S_{2}Y\right)$
@@
$\begin{cases}\widetilde{A}_{0}=A_{0}-A_{0}Q,\\ \widetilde{A}_{1}=A_{1}+A_{2}V_{2}S_{2}Y+XS_{1}^{-1}U_{1}{\widetilde{A}_{0}},\\ \widetilde{A}_{2}=A_{2}-PA_{2}.\end{cases}$
where $P=XU_{1}$, $Q=V_{2}Y$. Moreover, the eigenvalues of $\widetilde{A}(z)$ are those of $A(z)$ except for the eigenvalues of $S_{2}$, which are moved to 0, and the eigenvalues of $S_{1}^{-1}$, which are moved to infinity. In particular, $U_{1}\widetilde{A}_{2}=0$ and $\widetilde{A}_{0}V_{2}=0.$
@@
$A(z)=(I_{m}-zR)W(zI_{m}-G)$
where $R,G$ and $W$ are $m\times m$ matrices. Let $S_{1},S_{2},U_{1},V_{2}$, and $\widetilde{A}(z)$ be as in Theorem 2. Moreover, assume that $U_{1}R=S_{1}^{-1}U_{1}$ and $GV_{2}=V_{2}S_{2}$. Then the matrix polynomial $\widetilde{A}(z)$ can be factorized as
@@
where $\widetilde{G}=G-V_{2}S_{2}Y$ and $\widetilde{R}=R-XS_{1}^{-1}U_{1}$.
From the above theorem, it follows that the matrices $\widetilde{G}$ and $\widetilde{R}$ are the solutions of the QME $\widetilde{A}_{0}+\widetilde{A}_{1}X+\widetilde{A}_{2}X^{2}=0$ and $X^{2}\widetilde{A}_{0}+X\widetilde{A}_{1}+\widetilde{A}_{2}=0$, respectively, with eigenvalues the eigenvalues of $G$ and $R$, except for the ones coinciding with the eigenvalues of $S_{2}$ and $S_{1}^{-1}$, respectively, which are replaced by 0.
@@ 3 Deflation of the eigenvalues not lying on the unit circle
$A_{0}W_{G,1}+A_{1}W_{G,1}\Lambda_{G,1}+A_{1}W_{G,1}\Lambda_{G,1}^{2}=0,\penalty 10000\ \penalty 10000\ \penalty 10000\ \Lambda_{R,1}^{2}T_{R,1}A_{0}+\Lambda_{R,1}T_{R,1}A_{1}+T_{R,1}A_{2}=0.$ (11)
By applying Theorem 2 with $S_{1}^{-1}={\Lambda}_{R,1}$, $U_{1}=X^{\top}=T_{R,1}$, $S_{2}={\Lambda}_{G,1}$, $V_{2}=Y^{\top}=W_{G,1}$, we find that the function
$\widetilde{A}(z)=\left(I_{m}+zT_{R,1}^{\top}({\Lambda}_{R,1}^{-1}-zI_{q})^{-1}T_{R,1}\right)A(z)\left(I_{m}+W_{G,1}(zI_{q}-{\Lambda}_{G,1})^{-1}{\Lambda}_{G,1}W_{G,1}^{\top}\right)$
@@
Define $\tilde{G}=W_{G}^{\top}GW_{G}$, then
$\tilde{G}=\left[\begin{array}[]{ll}W_{G,2}^{\top}GW_{G,2}&W_{G,2}^{\top}GW_{G,1}\\ W_{G,1}^{\top}GW_{G,2}&W_{G,1}^{\top}GW_{G,1}\end{array}\right]=\left[\begin{array}[]{ll}W_{G,2}^{\top}GW_{G,2}&W_{G,2}^{\top}W_{G,1}\Lambda_{G,1}\\ W_{G,1}^{\top}GW_{G,2}&W_{G,1}^{\top}W_{G,1}\Lambda_{G,1}\end{array}\right]=\left[\begin{array}[]{ll}\bar{G}_{11}&0\\ \bar{G}_{21}&\Lambda_{G,1}\end{array}\right].$
Therefore,
$\begin{split}G&=W_{G}\tilde{G}W_{G}^{\top}\\ &=[W_{G,2}\penalty 10000\ W_{G,1}]\left[\begin{array}[]{ll}\bar{G}_{11}&0\\ \bar{G}_{21}&\Lambda_{G,1}\end{array}\right]\left[\begin{array}[]{ll}W_{G,2}^{\top}\\ W_{G,1}^{\top}\end{array}\right]\\ &=W_{G,2}\bar{G}_{11}W_{G,2}^{\top}+W_{G,1}\bar{G}_{21}W_{G,2}^{\top}+W_{G,1}\Lambda_{G,1}W_{G,1}^{\top}.\end{split}$
Define $\tilde{R}=T_{R}RT_{R}^{\top}$, then
$\tilde{R}=\left[\begin{array}[]{ll}T_{R,2}RT_{R,2}^{\top}&T_{R,2}RT_{R,1}^{\top}\\ T_{R,1}RT_{R,2}^{\top}&T_{R,1}RT_{R,1}^{\top}\end{array}\right]=\left[\begin{array}[]{ll}T_{R,2}RT_{R,2}^{\top}&T_{R,2}RT_{R,1}^{\top}\\ \Lambda_{R,1}T_{R,1}T_{R,2}^{\top}&\Lambda_{R,1}T_{R,1}T_{R,1}^{\top}\end{array}\right]=\left[\begin{array}[]{ll}\bar{R}_{11}&\bar{R}_{12}\\ 0&\Lambda_{R,1}\end{array}\right].$
Therefore,
$\begin{split}R&=T_{R}^{\top}\tilde{R}T_{R}\\ &=[T_{R,2}^{\top}\penalty 10000\ T_{R,1}^{\top}]\left[\begin{array}[]{ll}\bar{R}_{11}&\bar{R}_{12}\\ 0&\Lambda_{R,1}\end{array}\right]\left[\begin{array}[]{ll}T_{R,2}\\ T_{R,1}\end{array}\right]\\ &=T_{R,2}^{\top}\bar{R}_{11}T_{R,2}+T_{R,2}^{\top}\bar{R}_{12}T_{R,1}+T_{R,1}^{\top}\Lambda_{R,1}T_{R,1}.\end{split}$
@@ 4 Convergence properties of CR in the case of more than one eigenvalue on the unit circle
In this section, we derive some convergence properties of CR under the assumptions stated in Section 2.2. In this case, Theorem 1 does not hold, and CR generally does not converge if $\ell>1$. However, we will show that CR converges on a suitable subspace, and such a subspace can be used to derive the matrices $W_{G,1}$, $\Lambda_{G,1}$, $T_{R,1}$, and $\Lambda_{R,1}$ used in Section 3.
@@
Since the sequences $\{A_{1}^{(k)}\}_{k}$ and $\{A_{2}^{(k)}\}_{k}$ are bounded and $\rho(\Lambda_{G,1})=|\lambda_{m-\ell}|<1$, then we conclude that
@@
From the condition (7), as $k\rightarrow\infty$, $a_{k}(z)$ has exactly $m$-$\ell$ zeros that converge to 0, which implies that $\lim_{k\to\infty}{\operatorname{Null}(A_{0}^{(k)})}=m$-$\ell$. The properties for the sequence $\{A_{2}^{(k)}\}_{k}$ can be proved similarly, by multiplying the second equation in (6) to the left by $T_{R,1}$. ∎
@@
In particular, there exists a sequence of matrices $\{\widetilde{\Lambda}_{G,1}^{(k)}\}_{k}\in\mathbb{R}^{(m-l)\times(m-l)}$ such that
@@
is the SVD of $A_{2}^{(k)}$, then partition $U_{2}^{(k)}$ as $U_{2}^{(k)}=\begin{bmatrix}U_{2,1}^{(k)}\\ U_{2,2}^{(k)}\end{bmatrix}$, where $U_{2,1}^{(k)}\in\mathbb{R}^{\ell\times m}$, $U_{2,2}^{(k)}\in\mathbb{R}^{(m-\ell)\times m}$. As for the matrix $G$, there exists a sequence of matrices $\{\widetilde{\Lambda}_{R,1}^{(k)}\}_{k}\in\mathbb{R}^{(m-l)\times(m-l)}$ such that

### v2

@@ Abstract
organization=Department of Applied Mathematics, Lanzhou University of Technology,city=Lanzhou, postcode=730050, country=China
@@
organization=Dipartimento di Matematica, Università di Pisa,city=Pisa, postcode=56127, country=Italy
@@ 1 Introduction
A commonly used numerically stable approach for solving QMEs is to linearize them into associated generalized eigenvalue problems, which can then be efficiently solved using the QZ algorithm to compute all the solutions [21, 35, 15]. However, the computational cost limits its practicality for large-scale problems. Iterative methods, such as Newton’s method [21, 22], the Bernoulli iteration (BI) and its variants [21, 1], the fixed-point iterations [2, 29, 16, 13, 21, 7, 8], offer greater efficiency but may converge slowly or stagnate in nearly null-recurrent QBDs. An alternate to classical algorithms is the cyclic reduction (CR) algorithm. Originally developed by Golub and Hockney for block tridiagonal systems [10, 23], the CR algorithm was later extended to solve nonlinear matrix equations based on a functional representation [7, 30]. The central idea is to transform the nonlinear matrix equation into an equivalent semi-infinite, block Toeplitz, block tridiagonal or Hessenberg linear system and apply CR to solve this system. The comprehensive survey [9] provides a detailed account of the development and applications of CR.
@@ 2 Preliminaries and assumptions
Assume $a(z)$ not identically zero and denote its roots as $\lambda_{i}$, $i=1,\ldots,2m$, adding $s$ zeros at infinity if $a(z)$ has degree 2ms2m-s. Without loss of generality, assume that the roots are ordered according to their modulus as
@@ 2.1 Cyclic Reduction (CR)
Concerning convergence properties, we recall the following result, which follows from Theorems 9 and 10 of [9] (see also Theorem 7.6 in [7, Section 7.3]), and expresses the quadratic convergence:
@@
In the case where $\lambda_{m}=\lambda_{m+1}=1$, $|\lambda_{m-1}|<1$ and $|\lambda_{m+2}|>1$, under mild conditions, the convergence of CR turns to linear [17]. The convergence is restored to quadratic by applying a shifting technique to remove the eigenvalue 1 from the eigenvalues of $A(z)$ ([20], [7, Section 9.2]).
@@ 2.2 Assumptions
$A(z)=(I_{m}-zR)H(zI_{m}-G),$
where $H$ is a suitable nonsingular matrix; moreover, the matrix $A_{1}+A_{2}G$ is nosingular, and the matrices $G$ and $R$ are related by the equation (see [7, Section 3.3]):
@@
In this case, $\mu_{1},\ldots,\mu_{\ell}$ are the $\ell$-th roots of $1$. Moreover, $G$ and $R$ have nonnegative entries and, among all the possible solutions of (3) and (4), respectively, are the minimal ones according to the component-wise ordering (for details on these topics we refer the reader to [27], [7, Section 4.4], [14]).
@@ 2.3 Block shift-and-deflate technique
Denote by $\operatorname{rev}(A(z))$ the reversed quadratic matrix polynomial defined by
$\operatorname{rev}(A(z))=z^{2}A\!\left(z^{-1}\right)=A_{2}+zA_{1}+z^{2}A_{0},$
which is obtained by reversing the order of the coefficient matrices of $A(z)$. It is straightforward to verify that, for any $\lambda\in\mathbb{C}\setminus\{0\}$, $\lambda$ is an eigenvalue of $A(z)$ if and only if $\lambda^{-1}$ is an eigenvalue of $\operatorname{rev}(A(z))$. Under the convention that $1/\infty=0$ and $1/0=\infty$, the eigenvalue $\infty$ of $A(z)$ corresponds to the eigenvalue 0 of $\operatorname{rev}(A(z))$, and vice-versa.
The following result from [5] describes how a set of eigenvalues of $A(z)$ can be shifted to 0 via a right transformation.
Lemma 1.
[5, Section 3] Let $1\leq q<m$, $S_{2}\in\mathbb{C}^{q\times q}$, and $V_{2}\in\mathbb{C}^{m\times q}$ be full rank matrices such that
$A_{0}V_{2}+A_{1}V_{2}S_{2}+A_{2}V_{2}S_{2}^{2}=0.$
Let $Y\in\mathbb{C}^{q\times m}$ such that $YV_{2}=I_{q}$. Then the function
$A^{(R)}(z)=A(z)\left(I_{m}+V_{2}(zI_{q}-S_{2})^{-1}S_{2}Y\right)$
is a quadratic matrix polynomial, having $q$ eigenvalues at 0, and matrix coefficients
$\begin{cases}A^{(R)}_{0}=A_{0}-A_{0}Q,\\ A^{(R)}_{1}=A_{1}+A_{2}V_{2}S_{2}Y,\\ A^{(R)}_{2}=A_{2},\end{cases}$
where $Q=V_{2}Y$. Moreover, the eigenvalues of $A^{(R)}(z)$ are those of $A(z)$ except for the eigenvalues of $S_{2}$, which are moved to 0. In particular, $A^{(R)}_{0}V_{2}=0.$
An analogous result holds for the left eigenvectors. Although not explicitly stated in [5], the following lemma can be obtained by arguments similar to those used for the right transformation, allowing a set of eigenvalues to be shifted to 0.
Lemma 2.
Let $1\leq q<m$, $S_{1}\in\mathbb{C}^{q\times q}$, and $U_{1}\in\mathbb{C}^{q\times m}$ be full rank matrices such that
$U_{1}A_{0}+S_{1}U_{1}A_{1}+S_{1}^{2}U_{1}A_{2}=0.$
Let $X\in\mathbb{C}^{m\times q}$ such that $U_{1}X=I_{q}$. Then the function
$A^{(L)}(z)=\left(I_{m}+XS_{1}(zI_{q}-S_{1})^{-1}U_{1}\right)A(z)$
is a quadratic matrix polynomial, having $q$ eigenvalues at 0, and matrix coefficients
$\begin{cases}A^{(L)}_{0}=A_{0}-PA_{0},\\ A^{(L)}_{1}=A_{1}+XS_{1}U_{1}A_{2},\\ A^{(L)}_{2}=A_{2},\end{cases}$
where $P=XU_{1}$. Moreover, the eigenvalues of $A^{(L)}(z)$ are those of $A(z)$ except for the eigenvalues of $S_{1}$, which are moved to 0. In particular, $U_{1}A^{(L)}_{0}=0.$
The next lemma addresses the case where certain eigenvalues of $\operatorname{rev}(A(z))$ are shifted to 0, which is mathematically equivalent to shifting the corresponding eigenvalues of $A(z)$ to infinity.
Lemma 3.
Let $1\leq q<m$, $S_{1}\in\mathbb{C}^{q\times q}$ and let $U_{1}\in\mathbb{C}^{q\times m}$ be a full rank matrix such that
$S_{1}^{2}U_{1}A_{0}+S_{1}U_{1}A_{1}+U_{1}A_{2}=0.$
Let $X\in\mathbb{C}^{m\times q}$ such that $U_{1}X=I_{q}$. Then the function
$\widehat{A}^{(L)}(z)=\left(I_{m}+zXS_{1}(I_{q}-zS_{1})^{-1}U_{1}\right)A(z)$
is a quadratic matrix polynomial, having $q$ eigenvalues at infinity, and matrix coefficients
$\begin{cases}\widehat{A}^{(L)}_{0}=A_{0},\\ \widehat{A}^{(L)}_{1}=A_{1}+XS_{1}U_{1}A_{0},\\ \widehat{A}^{(L)}_{2}=A_{2}-PA_{2}.\end{cases}$
where $P=XU_{1}$. Moreover, the eigenvalues of $\operatorname{rev}(\widehat{A}^{(L)}(z))$ are those of $\operatorname{rev}(A(z))$ except for the eigenvalues of $S_{1}$, which are moved to 0. In particular, $U_{1}\widehat{A}^{(L)}_{2}=0.$
Proof.
Applying Lemma 2 to the reversed matrix polynomial $\operatorname{rev}(A(z))$, by shifting the eigenvalues of $S_{1}$ to 0, the matrix polynomial $\widehat{A}^{(L)}(z)$ is the reversed polynomial of
$\left(I_{m}+XS_{1}(zI_{q}-S_{1})^{-1}U_{1}\right)\operatorname{rev}(A(z)).$
The following theorem generalizes the previous results by providing a unified transformation that simultaneously shifts two sets of eigenvalues, associated with right and left eigenvectors, to 0 and infinity, respectively.
@@
Let $1\leq q<m$, $S_{1}\in\mathbb{C}^{q\times q}$, $S_{2}\in\mathbb{C}^{q\times q}$, and $U_{1}\in\mathbb{C}^{q\times m}$, $V_{2}\in\mathbb{C}^{m\times q}$ be full rank matrices such that
$S_{1}^{2}U_{1}A_{0}+S_{1}U_{1}A_{1}+U_{1}A_{2}=0,\penalty 10000\ \penalty 10000\ A_{0}V_{2}+A_{1}V_{2}S_{2}+A_{2}V_{2}S_{2}^{2}=0.$
@@
$\widetilde{A}(z)=\left(I_{m}+zXS_{1}(I_{q}-zS_{1})^{-1}U_{1}\right)A(z)\left(I_{m}+V_{2}(zI_{q}-S_{2})^{-1}S_{2}Y\right)$
@@
$\begin{cases}\widetilde{A}_{0}=A_{0}-A_{0}Q,\\ \widetilde{A}_{1}=A_{1}+A_{2}V_{2}S_{2}Y+XS_{1}U_{1}{\widetilde{A}_{0}},\\ \widetilde{A}_{2}=A_{2}-PA_{2}.\end{cases}$
where $P=XU_{1}$, $Q=V_{2}Y$. Moreover, among the eigenvalues of $A(z)$, those corresponding to $S_{2}$ are moved to 0. Similarly, among the eigenvalues of $\operatorname{rev}(A(z))$, those corresponding to $S_{1}$ are moved to 0. In particular, $\widetilde{A}_{0}V_{2}=0$ and $U_{1}\widetilde{A}_{2}=0.$
Proof.
The conclusions follow immediately by successively applying the right shift transformation from Lemma 1 and the left shift transformation from Lemma 3 to the matrix polynomial $A(z)$. ∎
@@
$A(z)=(I_{m}-zR)H(zI_{m}-G)$
where $R,G$ and $H$ are $m\times m$ matrices. Let $S_{1},S_{2},U_{1},V_{2}$, and $\widetilde{A}(z)$ be as in Theorem 2. Moreover, assume that $U_{1}R=S_{1}U_{1}$ and $GV_{2}=V_{2}S_{2}$. Then the matrix polynomial $\widetilde{A}(z)$ can be factorized as
@@
where $\widetilde{G}=G-V_{2}S_{2}Y$ and $\widetilde{R}=R-XS_{1}U_{1}$.
From the above theorem, it follows that the matrices $\widetilde{G}$ and $\widetilde{R}$ are the solutions of the QME $\widetilde{A}_{0}+\widetilde{A}_{1}X+\widetilde{A}_{2}X^{2}=0$ and $X^{2}\widetilde{A}_{0}+X\widetilde{A}_{1}+\widetilde{A}_{2}=0$, respectively, with eigenvalues the eigenvalues of $G$ and $R$, except for the ones coinciding with the eigenvalues of $S_{2}$ and $S_{1}$, respectively, which are replaced by 0.
@@ 3 Deflation of the eigenvalues not lying on the unit circle
$A_{0}W_{G,1}+A_{1}W_{G,1}\Lambda_{G,1}+A_{2}W_{G,1}\Lambda_{G,1}^{2}=0,\penalty 10000\ \penalty 10000\ \penalty 10000\ \Lambda_{R,1}^{2}T_{R,1}A_{0}+\Lambda_{R,1}T_{R,1}A_{1}+T_{R,1}A_{2}=0.$ (11)
By applying Theorem 2 with $S_{1}={\Lambda}_{R,1}$, $U_{1}=X^{\top}=T_{R,1}$, $S_{2}={\Lambda}_{G,1}$, $V_{2}=Y^{\top}=W_{G,1}$, we find that the function
$\widetilde{A}(z)=\left(I_{m}+zT_{R,1}^{\top}{\Lambda}_{R,1}(I_{q}-z{\Lambda}_{R,1})^{-1}T_{R,1}\right)A(z)\left(I_{m}+W_{G,1}(zI_{q}-{\Lambda}_{G,1})^{-1}{\Lambda}_{G,1}W_{G,1}^{\top}\right)$
@@
From (13) and (14), we obtain
$\begin{split}\widetilde{G}&=W_{G}\bar{G}W_{G}^{\top}\\ &=[W_{G,2}\penalty 10000\ W_{G,1}]\left[\begin{array}[]{ll}\bar{G}_{11}&0\\ \bar{G}_{21}&0\end{array}\right]\left[\begin{array}[]{ll}W_{G,2}^{\top}\\ W_{G,1}^{\top}\end{array}\right]\\ &=W_{G,2}\bar{G}_{11}W_{G,2}^{\top}+W_{G,1}\bar{G}_{21}W_{G,2}^{\top},\end{split}$
Subsequently, (24) is derived from $G=\widetilde{G}+W_{G,1}{\Lambda}_{G,1}W_{G,1}^{\top}$.
Similarly, the expression for (25) can also be obtained. ∎
@@ 4 Convergence properties of CR in the case of more than one eigenvalue on the unit circle
In this section, we derive some convergence properties of CR under the assumptions stated in Section 2.2. In this case, Theorem 1 is not applicable, and CR generally does not converge if $\ell>1$. However, we will show that CR converges on a suitable subspace, and such a subspace can be used to derive the matrices $W_{G,1}$, $\Lambda_{G,1}$, $T_{R,1}$, and $\Lambda_{R,1}$ used in Section 3.
@@
Since the sequences $\{A_{1}^{(k)}\}_{k=1}^{\infty}$ and $\{A_{2}^{(k)}\}_{k=1}^{\infty}$ are bounded and $\rho(\Lambda_{G,1})=|\lambda_{m-\ell}|<1$, then we conclude that
@@
From the condition (7), as $k\rightarrow\infty$, $a_{k}(z)$ has exactly $m$-$\ell$ zeros that converge to 0, which implies that $\lim_{k\to\infty}{\operatorname{Null}(A_{0}^{(k)})}=m$-$\ell$. The properties for the sequence $\{A_{2}^{(k)}\}_{k=1}^{\infty}$ can be proved similarly, by multiplying the second equation in (6) to the left by $T_{R,1}$. ∎
@@
In particular, there exists a sequence of matrices $\{\widetilde{\Lambda}_{G,1}^{(k)}\}_{k=1}^{\infty}\in\mathbb{R}^{(m-l)\times(m-l)}$ such that
@@
is the SVD of $A_{2}^{(k)}$, then partition $U_{2}^{(k)}$ as $U_{2}^{(k)}=\begin{bmatrix}U_{2,1}^{(k)}\\ U_{2,2}^{(k)}\end{bmatrix}$, where $U_{2,1}^{(k)}\in\mathbb{R}^{\ell\times m}$, $U_{2,2}^{(k)}\in\mathbb{R}^{(m-\ell)\times m}$. As for the matrix $G$, there exists a sequence of matrices $\{\widetilde{\Lambda}_{R,1}^{(k)}\}_{k=1}^{\infty}\in\mathbb{R}^{(m-l)\times(m-l)}$ such that